OpenStudy (anonymous):

A lathe machine produces machine parts whose lengths are normally distributed with a standard deviation of 0.20 mm. A sample of 10 parts is found to have a mean of 16.35 mm. Based on this sample, what is the 95% confidence interval for the population mean ?

OpenStudy (anonymous):

@rational

OpenStudy (rational):

start by finding the "margin of error"

OpenStudy (rational):

whats the multipilier(Z*) value for 95% confidence interval ? (look up in ur notes)

OpenStudy (anonymous):

.95

OpenStudy (rational):

there will be a multiplier value, look up ur notes once

OpenStudy (rational):

For 95% confidence interval, the multiplier value is `1.96` : 

OpenStudy (rational):

therefore margin of error is \[1.96*\frac{0.2}{\sqrt{10}} = ?\]

OpenStudy (anonymous):

I feel like I messed up

OpenStudy (rational):

messed up what

OpenStudy (anonymous):

.0153664 is what I got and that does not seem right to me

OpenStudy (rational):

work it again use ur calculator

OpenStudy (anonymous):

calc says .1214314622

OpenStudy (rational):

looks good, round it to 0.12 maybe

OpenStudy (rational):

so the margin of error is 0.12

OpenStudy (rational):

subtract that from mean, what do u get ?

OpenStudy (rational):

16.35-0.12 = ?

OpenStudy (anonymous):

16.23

OpenStudy (rational):

yes thats the lower bound of confidence interval

OpenStudy (rational):

for upper bound simply add 16.35 + 0.12 = ?

OpenStudy (anonymous):

16.47

OpenStudy (anonymous):

16.22-16.47

OpenStudy (rational):

so the confidence interval is \[\large (16.23,~ 16.47)\]

OpenStudy (anonymous):

Tyty